OpenStudy (anonymous):

perimeter of a quadrilateral is root 404 + root 500 + root 52 + root 52. I need to express this as 'a root b + c root d + e root 17

OpenStudy (anonymous):

please help!

OpenStudy (campbell_st):

well look at simplfiying a few \[\sqrt{404} = \sqrt{4 \times 101} = 2\sqrt{101}\] \[\sqrt{500} = \sqrt{100 \times 5} = 10\sqrt{5}\] and 52 = 4 x 13 so simplify each square root, then collect like terms and you will have a solution in the form you need

OpenStudy (anonymous):

cheers, but i was having trouble coming up with a like term to with the answer 'something root 17'

OpenStudy (campbell_st):

well based on the information posted, there isn't a sqrt17 value unless you have mistyped a value...